OpenStudy (anonymous):

Help Please.... When adding polynomials what do you do? how do you the expoenets? How do you divide and multiply polynomials. Please Help explain rules thanks

jimthompson5910 (jim_thompson5910):

you add like terms so x^2 and 3x^2 are like terms which means x^2 + 3x^2 = 4x^2 but x^2 and 3x^6 are NOT like terms, and you just leave it as x^2 + 3x^6

jimthompson5910 (jim_thompson5910):

you would do this for every term

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

Yes thank you what if it were in fraction form like this \frac{ 2y{ y(4y-1)} + \frac{ 5}{ 4y-1}

jimthompson5910 (jim_thompson5910):

you would have to make sure the denominators were equal before you could add the fractions

OpenStudy (anonymous):

Can you show me how please

jimthompson5910 (jim_thompson5910):

what's the LCD

OpenStudy (anonymous):

4y-1? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

y(4y-1) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

so each fraction needs to have an LCD of y(4y-1)

OpenStudy (anonymous):

4y^2-y @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no need to distribute

jimthompson5910 (jim_thompson5910):

you just need to make sure each denominator has y(4y-1) in it

OpenStudy (anonymous):

How do i make it that they both have y(4y-1)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

multiply top and bottom of the second fraction by y to get this

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ 2y}{ y(4y-1)} + \frac{ 5}{ 4y-1}\] \[\Large \frac{ 2y}{ y(4y-1)} + \frac{ 5y}{ y(4y-1)}\]

jimthompson5910 (jim_thompson5910):

now you can add the fractions

OpenStudy (anonymous):

\[\frac{ 7y }{ y(4y-1) }\] is that the answer? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

then reduce

OpenStudy (anonymous):

\[\frac{ 7y }{ 4y ^{2}-y }\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

notice a pair of terms will cancel

OpenStudy (anonymous):

The y? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

what's the final answer

OpenStudy (anonymous):

\[\frac{ 7 }{ 4y ^{2} }\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

it would just be \[\Large \frac{ 7}{ 4y-1 }\]

OpenStudy (anonymous):

oh ok and can i ask more questions please? i need you to revise my answers is that okay @jim_thompson5910

jimthompson5910 (jim_thompson5910):

how many questions are we talking about

OpenStudy (anonymous):

7 @jim_thompson5910

OpenStudy (anonymous):

If yourbusy thats ok you dont need to keep helping me youve done enough

jimthompson5910 (jim_thompson5910):

I can do a few, but idk about 7

OpenStudy (anonymous):

Ok thank you and are you sure? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah I'm sure

OpenStudy (anonymous):

Thank you=) \[(4a ^{4}b ^{3}+ 4a ^{3}b ^{2})+(5a ^{3}b ^{2}-4a ^{4}b ^{3})= 9a ^{3}b ^{2}\] is that right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

\[4xy(x+y)= 4x ^{2}y+4xy ^{2}\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

\[(x ^{3}+y ^{3})(x ^{3}-y ^{3})= x ^{6}-y ^{6}\]

jimthompson5910 (jim_thompson5910):

correct, I can do one more

OpenStudy (anonymous):

I need help on this one \[3x ^{7}+6x ^{6}-9x ^{5}+6x ^{4}\] i need to find gcf and i need to factor the expression. is the gcf 3x^7 and how do i factor the expression?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

the GCF is 3x^???

jimthompson5910 (jim_thompson5910):

replace ??? with the correct exponent (it's not 7)

OpenStudy (anonymous):

6? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

4? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

GCF is 3x^4

OpenStudy (anonymous):

why is it four? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

because x^7, x^6, etc aren't factors of x^4

jimthompson5910 (jim_thompson5910):

but x^4 is a factor of x^4, x^5, x^6, etc

OpenStudy (anonymous):

oh ok how do i facot out the expression?

jimthompson5910 (jim_thompson5910):

use the distributive property, just in reverse

OpenStudy (anonymous):

would it be \[3x ^{4}(x ^{3}+2x ^{2}-3x+2)\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, that's how far you can go

jimthompson5910 (jim_thompson5910):

you can't factor any further

OpenStudy (anonymous):

so thats the answer? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you so much for your help @jim_thompson5910